OpenStudy (anonymous):

Estimate the length of the radius, if the area of the circle is 750 mm. sq. A)19.78 B)23.87 C)9.36 D)15.45

OpenStudy (goformit100):

D)15.45

OpenStudy (anonymous):

HOW DID YOU GET THAT ?

OpenStudy (goformit100):

USE AREA OF CIRCLE=\[piR ^{2}\] WHERE r IS THE RADIUS

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

Thank You goformit100 (:

jhonyy9 (jhonyy9):

area = r^2*pi so than 750 =r^2 *pi us for pi =3,14 so than 750=r^2 * 3,14 r^2 =750/(3,14) =238,8 r=sqrt(238,8) = 15,45